OpenStudy (anonymous):

A fruit vendor has 28 pears and 16 oranges. What is the ratio of oranges to pears in simplest form?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

sorry @CandyCove

OpenStudy (flcon04):

Okay so for every 1 pear there is 1.75 oranges.

OpenStudy (flcon04):

So how many oranges are for 2 pears?

OpenStudy (anonymous):

thats the answer? @flcon04

OpenStudy (flcon04):

No

OpenStudy (flcon04):

Do you want the answer?

OpenStudy (flcon04):

Or do you want me to guide you?

OpenStudy (anonymous):

yes plaese @flcon04

OpenStudy (anonymous):

answer plz

OpenStudy (flcon04):

Okay so to find the answer what is 1.75 x 2?

OpenStudy (anonymous):

can u tell me the answer? @flcon04

OpenStudy (flcon04):

Its against code of conduct to just give out a answer so i cant but if you answer this the main answer will pop up.

OpenStudy (mathmale):

Your job is clearly stated: "What is the ratio of oranges to pears in simplest form? " Any questions about that? @dstella: Our Code of Conduct is clear in forbidding requests for direct answers. Please read this code if you have not done so already. Others and I would be glad to help you find the answer, but we are forbidden to hand out answers without your involvement and effort.

OpenStudy (mathmale):

http://openstudy.com/code-of-conduct

OpenStudy (anonymous):

ok

OpenStudy (flcon04):

So can you do 1.75 x 2?

OpenStudy (mathmale):

What is a "ratio?" Give an example. What is the ratio of the number of oranges to the number of pears?

OpenStudy (mathmale):

@flcon04 : "So can you do 1.75 x 2?" is neither a fair nor relevant question. How would dstella know where that came from and / or what it means? Better start with basics.

OpenStudy (flcon04):

Ok

563blackghost (563blackghost):

You would first setup the ration... So it states "A vendor machine has 28 pears and 16 oranges" that would make it a 28:16 ratio.... We now must find the greatest common factor they both have....so lets list the factors... 28:1, 2, 4, 7, 14, 28 16: 1,2,4,8,16 We now then choose which is the greatest common factor of both and we would then divide by that factor for both....in which we would then get the answer....

OpenStudy (whpalmer4):

28 pears and 16 oranges. Is there a simpler ratio? Think of having a group of friends where you wanted to divide up the fruit so that each friend got the same amount of fruit. How many friends could you have and make it come out evenly? If there are two of you, you could give each person 1/2 of the pears (28/2 = 14) and 1/2 of the oranges (16/2 = 8). So far, so good. But now two more friends show up. Can you divide up the fruit again so that everyone gets the same amount? Well, you and your first friend both have 14 pears and 8 oranges. Those are even numbers (can be divided by 2 without needing to make a fraction), so each of you could give half of each fruit to one of the two newcomers. That would give each of the 4 of you 14/2 = 7 pears and 8/2 = 4 oranges. Is it possible to divide up the fruit any further, if everyone is to have the same ratio of pears to oranges? If not, then the number of pears to oranges in the latest division is the simplest form. Note well that it may not be simply division by 2. What if you had 9 pears to 6 oranges? You can't divide that with 1 more friend, but you could divide it with 2 more friends (3 people total). In that case, you would have 3 pears and 2 oranges for each of the 3 people, and the simplest form of the ratio would be 3 pears to 2 oranges, or a ratio of pears:oranges of 3:2.